OpenStudy (anonymous):

what's the volume of a certain amount of gas at stp if this amount has occupied a volume of 250 cm^3 at room temperature and pressure 85 k.pa

OpenStudy (jfraser):

use the combined gas law\[\frac{P_1*V_1}{T_1} = \frac{P_2*V_2}{T_2}\]

OpenStudy (anonymous):

V=P'xV/P=85*250/0.84=25331.3CM^3 is this right?